If f is a function such that f(0)= 2 and f(1)= 3 and f(x+2)=f(x)-f(x+1) for every real x then f(5) equal to ?

Grade:12

1 Answers

Pratham Ashish
14 years ago

hi

ur question is quite simple, u just hav 2 use the given conditions till u get f(5)


f(x+2) =  f(x) - f(x+1)

putting x=0,

f(2) = f(0) - f(1)...................[x=0]

       = 2 - 3 = -1

f(3) = f(1) - f(2)....................[x=1]

     = 3 - (-1) = 4

f(4) = f(2) - f(3) ....................[x=2]

     = -1 - 4 = -5

f(5) = f(3) - f(4)....................[x=3]

      = 4- (-5) = 9